# Heterogeneous Treatment Effects and

### v1

@@ Abstract
The credibility revolution advances the use of research designs that permit identification and estimation of causal effects. However, understanding which mechanisms produce measured causal effects remains a challenge. A dominant current approach to the quantitative evaluation of mechanisms relies on the detection of heterogeneous treatment effects with respect to pre-treatment covariates. This paper develops a framework to understand when the existence of such heterogeneous treatment effects can support inferences about the activation of a mechanism. We show first that this design cannot provide evidence of mechanism activation without additional, generally implicit, assumptions. Further, even when these assumptions are satisfied, if a measured outcome is produced by a non-linear transformation of a directly-affected outcome of theoretical interest, heterogeneous treatment effects are not informative of mechanism activation. We provide novel guidance for interpretation and research design in light of these findings.
@@
The credibility revolution in empirical social science has motivated the largescale adoption of research designs that facilitate unbiased estimation of causal effects (Samii, 2016; Angrist and Pischke, 2010). Such unbiased estimates allow for valid measurement of the effect of a treatment on an outcome. However, they do not generally provide evidence about why or how the treatment affected the outcome. These questions of why or how are ultimately questions about the activation and influence of causal mechanisms. Understanding the mechanisms that produce causal effects is central to our ability to use empirical evidence to understand social phenomena (Slough and Tyson, 2023a).
@@
To examine the prevalence of the use of HTEs to learn about causal mechanisms, we survey the 2021 volumes of three leading journals in political science: the American Journal of Political Science (AJPS), the American Political Science Review (APSR), and the Journal of Politics (JoP). We first identify the subset of papers that analyze quantitative empirical data. We then report the proportion of those empirical papers that report HTEs or subgroup-specific treatment effects. Finally, we report the proportion of papers using HTEs that interpret these quantities as providing information about causal mechanisms(s). Table 1 shows that in each of the three leading journals, a majority of quantitative studies estimate HTEs. Moreover, conditional on reporting any HTEs, the vast majority of articles (82% across three journals) interpret these quantities as providing information about mechanisms. Collectively, these figures indicate that almost half (46%) of recent quantitative empirical articles in these journals use HTEs to assess mechanisms. Table A.1 further documents that the share of studies that use of HTE for mechanism detection is similar across all quantitative research designs/identification strategies in common usage.
@@
Existing concerns about HTEs have largely focused on the statistical properties of relevant estimators and hypothesis tests. In particular, interaction effects are known to have low statistical power (e.g., McClelland and Judd, 1993). Moreover, estimation of HTEs with respect to many (pre-treatment) covariates risks multiple comparisons problems (Gerber and Green, 2012; Lee and Shaikh, 2014; Fink, McConnell, and Vollmer, 2014). While these criticisms are important, they are distinct from theoretical questions about how the presence or absence of HTEs links to causal mechanisms. We take on this challenge by asking: under what conditions do HTEs provide evidence of mechanism activation?
@@
To answer this question, we develop a framework to formally link HTEs with respect to a covariate to the effect of a specific mechanism. To do so, we extend the workhorse causal mediation framework (Imai, Keele, and Tingley, 2010). Within our framework, a mechanism is an underlying process that influences experience in order to produce a (causal) effect if it is activated (Slough and Tyson, 2023a). A mediator, or mechanism representation, should thus be affected by treatment and have a non-zero (indirect) effect on the outcome if the mechanism is active for some unit. In order to use HTEs to detect mechanisms, empiricists rely on a measured moderator or pre-treatment covariate that is thought to predict the degree to which treatment activates a mechanism and/or the mechanism’s effect on an outcome of interest. Our results characterize the conditions under which heterogeneity in conditional average treatment effects (CATEs) with respect to a given moderator is a sufficient condition to show that the mechanism is active. A mechanism is active when its indirect effect is non-zero for some unit.
@@
We first present two classes of exclusion assumptions that link HTEs to a specified mechanism. These assumptions hold that the moderator of interest is excluded with respect to (1) the average indirect effect of other mechanisms and (2) the average direct effect of the treatment on the outcome. The intuition for these assumptions is straightforward. The use of HTEs to assess the activation of a given mechanism relies on the additive separability of the effect of that mechanism from the effects of other mechanisms. These exclusion assumptions are needed to ensure that additive separability is possible. In their absence, the relationship between HTEs and the indirect effect of a mechanism is unspecified. In this sense, these assumptions are implicitly invoked by current practice, but are not explicitly stated or defended.
@@
When these exclusion assumptions hold, what can we learn about a mechanism of interest from HTEs with respect to a given covariate? Our main results characterize what we can learn from the existence or non-existence of HTEs. We show that whether such learning is possible depends on the relationship between measured outcomes and the theory. We thus distinguish between outcome variables that are theorized to be directly affected by a mechanism versus (causally) subsequent outcomes that are indirectly affected.
@@
First, consider the case in which a measured outcome is directly affected by the mechanism of interest. We show that when both exclusion assumptions hold, the existence of HTEs provides evidence of mechanism activation. This broadly conforms to current practices in applied research, albeit while making the underlying assumptions explicit. However, when there do not exist HTEs, there are two possible explanations: (1) the indirect effect of a mechanism is not moderated by any covariate; and/or (2) the theorized relationship between the covariate and the indirect effect of the mechanism was misspecified. Neither possibility distinguishes between an active or inactive mechanism. This means that a lack of HTEs cannot “rule out” a mechanism by showing that it is inactive. This finding contradicts standard interpretations of (the absence of) HTEs in the applied literature.
@@
We further consider the common case in which we observe outcomes that are indirectly affected by a mechanism. For example, in a political economy model, a mechanism may change an actor’s utility, but researchers measure only the actor’s discrete choice/behavior. In political psychology, a treatment may affect a subject’s (latent) attitudes, but researchers measure their survey response on a Likert scale. When an outcome is generated by a non-linear mapping of the directly-affected outcome (as in both of these examples), the relationship between HTEs and mechanisms changes. In this case, the existence of HTEs with respect to a covariate no longer provides evidence of mechanism activation, even when both exclusion assumptions hold. The logic for this result is straightforward: the non-linear mapping of the mechanism’s influence into an observed outcome breaks the additive separability of the indirect effect of interest from other indirect and direct effects, which undermines our ability to link HTEs to the indirect effect of a mechanism.
@@
This paper makes three principal contributions. First, we contribute to literature on varying uses of HTEs. Our focus, in line with most current empirical applications, is on the use of HTEs to learn about the mechanisms that produce treatment effects. It is crucial to distinguish this of HTEs from the use of HTEs for extrapolation, prediction, or targeting of treatment. The latter class of concerns—extrapolation, prediction, and targeting—that are more dominant in the recent methodological literature (Egami and Hartman, 2022; Devaux and Egami, 2022; Kitagawa and Tetenov, 2018; Athey and Wager, 2021). Whereas these applications are enhanced by the use of machine learning methods for the detection of heterogeneity (e.g., Grimmer, Messing, and Westwood, 2017; Athey, Tibshirani, and Wager, 2019), we show that using HTEs to learn about mechanisms instead relies on a deductive theoretical mapping of covariates to mechanisms, which is unlikely to be advanced by developments in machine learning.
@@
Second, we expand a growing literature on the theoretical implications of empirical models (TIEM) (Bueno de Mesquita and Tyson, 2020; Ashworth, Berry, and Bueno de Mesquita, 2021; Ashworth, Berry, and de Mesquita, 2023; Abramson, Koçak, and Magazinnik, 2022; Slough, 2022). We make two central interventions to this literature. First, our framework makes explicit links between a causal mediation framework that is used more prominently by empiricists and (formal) theoretical models. Second, we introduce questions about how measured outcomes relate to theoretical constructs by distinguishing between directly- and indirectly-affected outcomes. While measurement is central to recent TIEM work on evidence accumulation in a cross-study environment (Slough and Tyson, 2023a, b, c), it has not been widely explored in the single-study environment.
@@
Finally, we provide practical guidance for empirical researchers who want to learn about which mechanisms generate observed effects. Our assumptions and results reveal a minimal set of attributes of an applied theory that can support the use of HTEs to learn about mechanisms. Two of these attributes, the relationships between (1) a covariate of interest and other mechanisms and (2) measured outcomes and theoretical objects of interest, are generally not discussed in applied work. Second, we show how interpretation of HTEs can be improved, returning to the statistical problems that are well known in this literature. Third, we discuss how our analysis can be used to inform prospective research design. Last, we show that in order to infer mechanism activation from HTEs in the case of indirectly-observed outcomes, stronger theoretical assumptions about the mapping from directly-observed to indirectly-observed outcomes are necessary. Collectively, these suggestions allow practitioners to accurately use—or, when indicated, avoid—HTEs as a quantitative test of mechanisms.
@@ 1 Current Practice
As reported in Table 1, 82% of the articles that report HTEs interpret these quantities as tests of a mechanism. Two interpretations of HTEs are common. First, the presence of HTEs with respect to a specific covariate provides evidence that a mechanism is active. For example, Haim, Ravanilla, and Sexton (2021) report the results of a field experiment in a conflict-affect region of the Philippines that randomized provision of a program that connected villages to state services and sought to increase village leaders’ trust in the state. In the context of the COVID-19 pandemic, they show that the program increased village leaders’ probability of reporting COVID risk information to a government task force by $\approx 10$ percentage points. Haim, Ravanilla, and Sexton (2021) argue that the program increased reporting by increasing beliefs in government competence. They report HTEs that indicate the effect on reporting was significantly larger in communities in which village leaders initially believed that the government “[did] not have capacity to meet needs.” Here, the presence of HTEs with respect to perceived capacity is thought to provide evidence that leaders assigned to treatment updated their beliefs about government capacity (the mechanism), thereby increasing their willingness to report risk information to the government.
Second, the absence of HTEs with respect to a given covariate is frequently used to “rule out” the activation of a possible mechanism (often called an alternative explanation). For example, Moscowitz (2021) argues that the proportion of in-state residents in one’s local media market increases rates of voter political knowledge and split-ticket voting in the US, countering trends toward the nationalization of politics. The paper provides evidence that news coverage is the mechanism that drives this effect. However, it also seeks to rule out an alternative mechanism: campaign advertising. The logic for this alternative mechanism that holds that more in-state residents leads to more campaign advertising about in-state candidates, which in turn increases voter knowledge. Since advertisements air when during election season (when incumbents contest re-election), Moscowitz (2021) codes an indicator that takes the value of “1” when the survey was fielded during an election season and “0” otherwise. The effect of the share of in-state residents in a local media market does not change during election seasons. This lack of heterogeneity is used to provide evidence against the advertisement-based mechanism.
@@ 2 Motivating Example: Exogenous Shocks and Voting
Consider a large class of natural experiments on the effect of some exogenous shock, denoted $\omega$, on voter beliefs and behavior in a democracy. [^3] One could alternatively conduct a survey or field experiment in which researchers provide information on an exogenous shock that is otherwise imperfectly observed by voters. This shock could be a natural disaster (e.g. Healy and Malhotra, 2010; Achen and Bartels, 2017), a disease outbreak (e.g., Baccini, Brodeur, and Weymouth, 2021), an economic crisis (e.g. Wolfers, 2002), or even a seemingly-irrelevant event (e.g. Healy, Malhotra, and Mo, 2010). In order to characterize the effect of the shock on voter beliefs and behavior, we adapt a formal model by Ashworth, Bueno de Mesquita, and Friedenberg (2018).
@@
Each voter’s utility from a politician depends on the politician’s type, $\theta$, and a valence shock for the incumbent, $v_{i}$. Here, valence captures any attribute of the incumbent that does not depend on their type, including but not limited to bias toward a candidate or ideological closeness. Politician type and valence are additively separable, such that voter utility is given by:
@@ 2.1 From Theory to Empirical Research Design
Mapping this model onto the empirical research design, we will assume that $\omega\in\{\omega^{\prime},\omega^{\prime\prime}\}$ denotes a binary treatment, where $\omega^{\prime}$ indicates no exposure to the shock and $\omega^{\prime\prime}$ denotes exposure to the shock. The first outcome, $y_{1}(\omega)$, measures voters’ expected utility from the incumbent. It is obviously difficult and rare to measure utility directly, though one could, in principle, elicit willingness-to-pay. A voter’s expected utility from a vote for the incumbent is given by:
@@
Our model can be represented as a directed acyclic graph (DAG), as depicted in Figure 1. This representation clarifies a number of assumptions of the model. First, the mechanism through which the shock affects voter preferences and behavior is through voter learning. This is evident because the only path from the shock ($\omega$) to the outcomes $y_{1}(\omega)$ and $y_{2}(\omega)$ passes through the voter’s posterior beliefs about the incumbent’s type. Voter learning is moderated by the voter’s prior belief about the incumbent, $\pi_{i}(\overline{\theta})$. Throughout this paper, we will indicate causal moderation through the arrow pointing to a path rather than a node, following the suggestion of Weinberg (2007). [^6] We could define an extra node to indicate this interaction. We do not do so in the interest of parsimony. Note that the representation of “interaction” effects in DAGs is not standardized (Nilsson et al., 2021). Because our assumptions encode more structure (e.g., excluding the possibility of certain moderation effects), we need to encode additional structure in graphical representations than is standard. For the purposes of exposition, the mediator, posterior beliefs, is unobserved to researchers. We will assume, however, that researchers have a measure of $\pi_{i}$ and $v_{i}$ from a baseline survey.
@@
Under our model of voter updating and behavior, do heterogeneous treatment effects provide evidence that the relevant mechanism is voter learning? Ex-ante, the empiricists do not know that learning is the active (or operative) mechanism. To learn about mechanisms, many researchers will estimate HTEs, typically pointing to heterogeneity as evidence of mechanism activation. When is this approach valid? When does it yield invalid substantive inferences about mechanisms? To develop intuitions, we evaluate four HTE combinations using moderators $\textbf{X}=\{\Pi,V\}$, where $\Pi$ is the set of all possible values of $\pi_{i}$ and $V$ is the set of all possible values of $v_{i}$, and outcomes $y\in\{y_{1},y_{2}\}$. Remark 1 shows that for the outcome measuring a voter’s expected utility from a vote for the incumbent ($y_{1})$, detecting heterogeneity in CATEs correctly provides evidence that the mechanism is voter learning about the incumbent’s type.
@@
Why do we see HTEs in $v$ for vote choice? Recall that each voter votes for the incumbent if and only if $\beta(\overline{\theta}|\pi_{i},\omega)+v_{i}-\pi^{C}\geq 0$. But this binary choice means that voter beliefs and valence are no longer additively separable with respect to vote choice. Consequently, with a sufficiently large sample size (and thus sufficient statistical power), researchers are apt to detect HTEs even when a mechanism is not active. Using standard interpretations of these tests, this leads to Type-I errors in our inferences about mechanism activation.

### v2

@@ Abstract
The credibility revolution advances the use of research designs that permit identification and estimation of causal effects. However, understanding which mechanisms produce measured causal effects remains a challenge. A dominant current approach to the quantitative evaluation of mechanisms relies on the detection of heterogeneous treatment effects with respect to pre-treatment covariates. This paper develops a framework to understand when the existence of such heterogeneous treatment effects can support inferences about the activation of a mechanism. We show first that this design cannot provide evidence of mechanism activation without an additional, generally implicit, assumption. Further, even when this assumption is satisfied, if a measured outcome is produced by a non-linear transformation of a directly-affected outcome of theoretical interest, heterogeneous treatment effects are not informative of mechanism activation. We provide novel guidance for interpretation and research design in light of these findings.
@@
The credibility revolution in empirical social science has motivated the largescale adoption of research designs that facilitate unbiased estimation of causal effects (Samii, 2016; Angrist and Pischke, 2010). Such unbiased estimates allow for valid measurement of the effect of a treatment on an outcome. However, they do not generally provide evidence about why or how the treatment affected the outcome. These questions of why or how are ultimately questions about the activation and influence of causal mechanisms. Understanding the mechanisms that produce causal effects is central to our ability to use empirical evidence to understand social phenomena (Slough and Tyson, 2024a).
@@
To examine the prevalence of the use of HTEs to learn about causal mechanisms, we survey the 2021 volumes of three leading journals in political science: the American Journal of Political Science (AJPS), the American Political Science Review (APSR), and the Journal of Politics (JoP). We first identify the subset of papers that analyze quantitative empirical data. We then report the proportion of those empirical papers that report HTEs or subgroup-specific treatment effects. Finally, we calculate the proportion of papers using HTEs that interpret these quantities as providing information about causal mechanisms(s). Table 1 shows that in each of the three leading journals, a majority of quantitative studies estimate HTEs. Moreover, conditional on reporting any HTEs, the vast majority of articles (82% across three journals) interpret these quantities as providing information about mechanisms. Collectively, these figures indicate that almost half (46%) of recent quantitative empirical articles in these journals use HTEs to assess mechanisms. Table A.1 further documents that the share of studies that use of HTE for mechanism detection is similar across all quantitative research designs/identification strategies in common usage.
@@
Existing concerns about HTEs have largely focused on the statistical properties of relevant estimators and hypothesis tests. In particular, interaction terms in regression—the most common estimators of HTEs—are known to have low statistical power (e.g., McClelland and Judd, 1993). Moreover, estimation of HTEs with respect to many (pre-treatment) covariates risks multiple comparisons problems (Gerber and Green, 2012; Lee and Shaikh, 2014; Fink, McConnell, and Vollmer, 2014). While these criticisms are important, they are distinct from theoretical questions about how the presence or absence of HTEs links to causal mechanisms. We take on this challenge by asking: under what conditions do HTEs provide evidence of mechanism activation?
@@
To answer this question, we develop a framework to formally link HTEs with respect to a covariate to the effect of a specific mechanism. To do so, we extend the workhorse causal mediation framework (Imai, Keele, and Tingley, 2010). Within our framework, a mechanism is an underlying process that influences experience in order to produce a (causal) effect if it is activated (Slough and Tyson, 2024a). A mediator, or mechanism representation, should thus be affected by treatment and have a non-zero (indirect) effect on the outcome if the mechanism is active for some unit. In order to use HTEs to detect mechanisms, empiricists rely on a measured moderator or pre-treatment covariate that is thought to predict the degree to which treatment activates a mechanism and/or the mechanism’s effect on an outcome of interest. Our results characterize the conditions under which heterogeneity in conditional average treatment effects (CATEs) with respect to a given moderator is a sufficient condition to show that the mechanism is active. A mechanism is active when its indirect effect is non-zero for some unit.
@@
We first present the exclusion assumption that links HTEs to a specified mechanism. The assumption holds that the moderator of interest is excluded with respect to other (average) effects of the treatment on the outcome. The intuition for the assumption is straightforward. The use of HTEs to assess the activation of a given mechanism relies on the additive separability of the effect of that mechanism from the effects of other mechanisms. The exclusion assumption is needed to ensure that additive separability is possible. In the absence of such an assumption, the relationship between HTEs and the indirect effect of a mechanism is unspecified. In this sense, this assumption is implicitly invoked by current practice, but is not explicitly stated or defended.
@@
When the exclusion assumption holds, what can we learn about a mechanism of interest from HTEs with respect to a given covariate? Our main results characterize what we can learn from the existence or non-existence of HTEs. We show that whether such learning is possible depends on the relationship between measured outcomes and the theory. We thus distinguish between outcome variables that are theorized to be directly affected by a mechanism versus (causally) subsequent outcomes that are indirectly affected.
@@
First, consider the case in which a measured outcome is directly affected by the mechanism of interest. We show that when the exclusion assumption holds, the existence of HTEs provides evidence of mechanism activation. This broadly conforms to current practices in applied research, albeit while making the underlying assumption explicit. However, when there do not exist HTEs, there are two possible explanations: (1) the indirect effect of a mechanism is not moderated by any covariate; and/or (2) the theorized relationship between the covariate and the indirect effect of the mechanism was misspecified. Neither possibility distinguishes between an active or inactive mechanism. This means that a lack of HTEs cannot “rule out” a mechanism by showing that it is inactive. This finding contradicts standard interpretations of (the absence of) HTEs in the applied literature.
@@
We further consider the common case in which we observe outcomes that are indirectly affected by a mechanism. For example, in a political economy model, a mechanism may change an actor’s utility, but researchers measure only the actor’s discrete choice/behavior. In political psychology, a treatment may affect a subject’s (latent) attitudes, but researchers measure their survey response on a Likert scale. When an outcome is generated by a non-linear mapping of the directly-affected outcome (as in both of these examples), the relationship between HTEs and mechanisms changes. In this case, the existence of HTEs with respect to a covariate no longer provides evidence of mechanism activation, even when the exclusion assumption holds. The logic for this result is straightforward: the non-linear mapping of the mechanism’s influence into an observed outcome breaks the additive separability of the indirect effect of interest from other indirect and direct effects, which undermines our ability to link HTEs to the indirect effect of a mechanism.
@@
This paper makes three principal contributions. First, our framework clarifies distinctions between varying uses of HTEs. Our focus, in line with most current empirical applications, is on the use of HTEs to learn about the mechanisms that produce treatment effects. This use of HTEs is distinct from important recent methodological contributions that emphasize the use of HTEs for extrapolation, prediction, or targeting of treatments (e.g., Egami and Hartman, 2022; Huang, 2022; Devaux and Egami, 2022; Kitagawa and Tetenov, 2018; Athey and Wager, 2021). Whereas these applications are facilitated by the application of machine learning methods (e.g., Grimmer, Messing, and Westwood, 2017; Chernozhukov et al., 2018; Athey, Tibshirani, and Wager, 2019), we show that using HTEs to learn about mechanisms instead relies on a deductive theoretical mapping of covariates to mechanisms, which is unlikely to be advanced by more widespread use of machine learning. Clarity about this distinction will improve interpretation of HTEs in applied work. It will further contribute to the development of methods for the multiple uses of HTEs.
@@
Second, we expand a growing literature on the theoretical implications of empirical models (TIEM) (Bueno de Mesquita and Tyson, 2020; Ashworth, Berry, and Bueno de Mesquita, 2021; Ashworth, Berry, and de Mesquita, 2023; Abramson, Koçak, and Magazinnik, 2022; Slough, 2023). We make two central interventions to this literature. First, our framework makes explicit links between a causal mediation framework that is used more prominently by empiricists and (formal) theoretical models. Second, we introduce questions about how measured outcomes relate to theoretical constructs by distinguishing between directly- and indirectly-affected outcomes. While measurement is central to recent TIEM work on evidence accumulation in a cross-study environment (Slough and Tyson, 2023, 2024a, 2024b), it has not been widely explored in the single-study environment.
@@
Finally, we provide practical guidance for empirical researchers who want to learn about which mechanisms generate observed effects. Our assumption and results reveal a minimal set of attributes of an applied theory that can support the use of HTEs to learn about mechanisms. Two of these attributes, the relationships between (1) a covariate of interest and other mechanisms and (2) measured outcomes and theoretical objects of interest, are generally not discussed in applied work. Second, we show how interpretation of HTEs can be improved, returning to the statistical problems that are well known in this literature. Third, we discuss how our analysis can be used to inform prospective research design. Finally, we show that in order to infer mechanism activation from HTEs in the case of indirectly-observed outcomes, stronger theoretical assumptions about the mapping from directly-observed to indirectly-observed outcomes are necessary. Collectively, these suggestions allow practitioners to accurately use—or, when indicated, avoid—HTEs as a quantitative test of mechanisms.
@@ 1 Current Practice
As reported in Table 1, 82% of the articles that report HTEs interpret these quantities as tests of a mechanism. Two interpretations of HTEs are common. First, the presence of HTEs with respect to a specific covariate provides evidence that a mechanism is active. For example, Haim, Ravanilla, and Sexton (2021) report the results of a field experiment in a conflict-affect region of the Philippines that randomized provision of a program that connected villages to state services and sought to increase village leaders’ trust in the state. In the context of the COVID-19 pandemic, they show that the program increased village leaders’ probability of reporting COVID risk information to a government task force by $\approx 10$ percentage points. Haim, Ravanilla, and Sexton (2021) argue that the program increased reporting by increasing beliefs in government competence. They report HTEs that indicate the effect on reporting was significantly larger in communities in which village leaders initially believed that the government “[did] not have capacity to meet needs.” These HTEs with respect to perceived capacity suggest that leaders assigned to treatment updated their beliefs about government capacity (the mechanism), thereby increasing their willingness to report risk information to the government.
Second, the absence of HTEs with respect to a given covariate is frequently used to “rule out” the activation of a possible mechanism (often called an alternative explanation). For example, Moscowitz (2021) argues that the proportion of in-state residents in one’s local media market increases rates of voter political knowledge and split-ticket voting in the US, countering trends toward the nationalization of politics. The paper provides evidence that news coverage is the mechanism that drives this effect. However, it also seeks to rule out an alternative mechanism: campaign advertising. The logic for this alternative mechanism holds that more in-state residents leads to more campaign advertising about in-state candidates, which in turn increases voter knowledge. Since advertisements air when during election season (when incumbents contest re-election), if the advertising mechanism were operative, we would expect the effect of in-state residents on voter knowledge to be larger during election seasons than other times in the electoral cycle. Moscowitz (2021) finds no such heterogeneity. He uses this finding is used to provide evidence against the advertisement-based mechanism.
@@ 2 Motivating Example: Exogenous Shocks and Voting
Consider a large class of natural experiments on the effect of some exogenous shock, denoted $\omega$, on voter beliefs and behavior in a democracy. [^3] One could alternatively conduct a survey or field experiment in which researchers provide information on an exogenous shock that is otherwise imperfectly observed by voters. This shock could be a natural disaster (e.g., Healy and Malhotra, 2010; Achen and Bartels, 2017), a disease outbreak (e.g., Baccini, Brodeur, and Weymouth, 2021), an economic crisis (e.g., Wolfers, 2002), or even a seemingly-irrelevant event (e.g., Healy, Malhotra, and Mo, 2010). In order to characterize the effect of the shock on voter beliefs and behavior, we adapt a formal model by Ashworth, Bueno de Mesquita, and Friedenberg (2018).
@@
Each voter’s utility from a politician depends on the politician’s type, $\theta$, and a valence shock for the incumbent, $v_{i}$. Here, valence captures any attribute of the incumbent that does not depend on their type, including, but not limited to, bias toward a candidate or ideological proximity. Politician type and valence are additively separable, such that voter utility is given by:
@@ 2.1 From Theory to Empirical Research Design
Mapping this model onto the empirical research design, we will assume that $\omega\in\{\omega^{\prime},\omega^{\prime\prime}\}$ denotes a binary treatment, where $\omega^{\prime}$ indicates no exposure to the (adverse) shock and $\omega^{\prime\prime}$ denotes exposure to the shock. The first outcome, $y_{1}(\omega)$, measures voters’ expected utility from the incumbent. It is obviously difficult and rare to measure utility directly, though one could, in principle, elicit willingness-to-pay. A voter’s expected utility from a vote for the incumbent is given by:
@@
Our model can be represented as a directed acyclic graph (DAG), as depicted in Figure 1. This representation clarifies a number of assumptions of the model. First, the mechanism through which the shock affects voter preferences and behavior is voter learning. This is evident because the only path from the shock ($\omega$) to the outcomes $y_{1}(\omega)$ and $y_{2}(\omega)$ passes through the voters’ posterior beliefs about the incumbent’s type. Voter learning is moderated by voters’ prior beliefs about the incumbent, $\pi_{i}(\overline{\theta})$. Throughout this paper, we will indicate causal moderation through the arrow pointing to a path rather than a node, following the suggestion of Weinberg (2007). [^6] We could define an extra node to indicate this interaction. We do not do so in the interest of parsimony. Note that the representation of “interaction” effects in DAGs is not standardized (Nilsson et al., 2021). Because our theoretical model encodes more structure (e.g., excluding the possibility of certain causal moderation effects), this notational convention encodes more structure in graphical representations than is standard. For the purposes of exposition, the mediator, posterior beliefs, is unobserved to researchers. We will assume, however, that researchers have a measure of $\pi_{i}$ and $v_{i}$ from a baseline survey.
@@
Under this model of voter updating and behavior, do heterogeneous treatment effects provide evidence that the relevant mechanism is voter learning? Ex-ante, the empiricists do not know that learning is the active (or operative) mechanism. To learn about mechanisms, many researchers will estimate HTEs, typically pointing to heterogeneity as evidence of mechanism activation. When is this approach valid? When does it yield invalid substantive inferences about mechanisms? To develop intuitions, we evaluate four HTE combinations using moderators $\textbf{X}=\{\Pi,V\}$, where $\Pi$ is the set of all possible values of $\pi_{i}$ and $V$ is the set of all possible values of $v_{i}$, and outcomes $y\in\{y_{1},y_{2}\}$. Remark 1 shows that for the outcome measuring a voter’s expected utility from a vote for the incumbent ($y_{1})$, heterogeneity in CATEs correctly provides evidence that the mechanism is voter learning about the incumbent’s type.
@@
Why do we see HTEs in $v$ for vote choice? Recall that each voter votes for the incumbent if and only if $\beta(\overline{\theta}|\pi_{i},\omega)+v_{i}-\pi^{C}\geq 0$. But this binary choice means that voter beliefs and valence are no longer additively separable with respect to vote choice. Consequently, with a sufficiently large sample size (and thus sufficient statistical power), researchers are apt to detect HTEs even when a mechanism is not active. The standard interpretation of this test would lead to a Type-I error in our inference about the activation of the valence mechanism.